(1 point) Solve the heat problem U = Uxx + sin(x) – 2 sin(2x), 0 < x < 1, u(0,t) = 0, u(,t) = 0 u(x,0) = 0 u(x,t) = sin(x)(1-My answers are wrong, please help
Answer #1

Ginen PDE - U ocx 2 ТТ = Uxxi + sinal – 2sin 2x is 4@, t) = u(, t) =0 46%,0)=0 taking Laplace troms form in PDE L{4} = Laseryū (x, t) = Cof. + Po I I sina - 2 sfman q e + G2 e t ū Bext) = s(8+4) + sistu a Sestu • 460,t) 30 ū (0, +) =0 os from . 0 = +lim uret) = 8 fun tim (-et) - I sinate lim (- ett) do tro lim uerst) = sina C1-0%) I sin 22 (1-2) I do = símu (1-0) - € sinen